# EDINET-Bench: Evaluating LLMs on Complex Financial Tasks using Japanese Financial Statements

### v1

@@ 4 Evaluation
We evaluate the zero-shot performance of current state-of-the-art LLMs, along with classical baselines, on EDINET-Bench.
@@ 4.1 Evaluation Setup
Baseline models
We evaluate seven frontier LLMs in a zero-shot setting, including GPT-4o (gpt-4o-2024-11-20) and o4-mini (o4-mini-2025-04-16) from OpenAI, Claude 3.5 Haiku (claude-3-5-haiku-20241022) and Claude 3.5 Sonnet (claude-3-5-sonnet-20241022) and Claude 3.7 Sonnet (claude-3-7-sonnet-20250219) from Anthropic, and DeepSeek-V3999https://openrouter.ai/deepseek/deepseek-chat and DeepSeek-R1101010https://openrouter.ai/deepseek/deepseek-r1 from DeepSeek. For inference, the temperature parameter is set to 0.0 and the maximum token length for generation is set to 4096. For GPT-4o, the random seed is fixed at 0. We use the test split for evaluation. We use the test split for evaluation and perform three runs for LLMs, as predictions with these models are not fully deterministic even with the temperature set to 0.0. We also use following baseline models:
  • Logistic regression: We use this model for accounting fraud detection and earnings forecasting tasks. We train this model on the training split and evaluate it on the test split. We use scikit-learn (Pedregosa et al., 2011) for the implementation, employing its default hyperparameter settings. To avoid excessive input attributes, only the Summary section of the annual report is used as the input feature.
  • Naive prediction: This model is used solely for earnings forecasting, predicting this year’s change will follow last year’s trend.
Prompt
We use the following system prompt: “You are a financial analyst.” For each task, we use the instruction prompt shown in Appendix A. For the binary classification tasks of accounting fraud detection and earnings forecasting, we prompt the models to output a prediction (0 or 1), a probability value (between 0 to 1), and the reasoning behind their judgment based on information from the securities reports. For the multi-class industry prediction task, we prompt the models to output the industry class and their reasoning for the classification.
Data processing
We extract the following items from the TSV files using edinet2dataset for prediction: Balance Sheet (BS), Cash Flow (CF), Profit and Loss (PL), Consolidated Summary (Summary), and Text Block (Text). These items are concatenated, appended to the end of the instruction prompt, and fed into the models in a zero-shot manner. To assess how information quantity affects LLM performance, we tested three configurations: Summary only, BS+CF+PL+Summary, and all items (BS+CF+PL+Summary+Text). Note that BS, CF, PL, and Summary consist of tabular data and do not contain company names or specific year information, whereas Text includes unstructured textual content that contains both company names and year information.
Cost and processing time analysis
@@
[…]
@@
[…]
Logistic Summary 0.68 0.17 0.56 0.05
Naive Prediction 0.44 -0.12
@@ 4.2 Results
Table 6 shows the performance of each model on each task in EDINET-Bench.
Accounting fraud detection
In the accounting fraud detection task, all LLMs faced challenges, with even the state-of-the-art Claude 3.5 Sonnet only slightly surpassing logistic regression in terms of ROC-AUC. Notably, most models demonstrated substantial improvements when text information was incorporated, indicating effective utilization of textual data for fraud detection.
Logistic regression, despite its simplicity, achieved surprisingly strong results, only slightly behind the best model/setup (Claude 3.5 that uses Text) in fraud detection. This highlights the substantial room for LLM improvement.
We take a closer look at the feature importance of logistic regression, shown in Table 7. This reveals that total comprehensive income and total assets were influential factors in non-fraud predictions, suggesting a bias towards predicting larger companies as non-fraud and smaller ones as fraud. Furthermore, the confusion matrix of Claude 3.7 Sonnet, shown in Figure 4, illustrates how incorporating textual information improved Sonnet’s accuracy in classifying both fraud and non-fraud cases.
Earnings forecasting
For the earnings forecasting task, many LLMs struggled to deliver precise predictions. Even Claude 3.7 Sonnet achieved only a ROC-AUC of 0.61, aligning with the low forecasting performance reported by Xie et al. (2024). This limitation likely stems from the challenge of predicting next-year earnings based solely on the current securities report. Unlike the fraud detection task, incorporating text information did not enhance performance in earnings forecasting, suggesting a weaker reliance on textual factors.
@@
Figure 3: Confusion matrix of logistic regression on accounting fraud detection.
Accounting fraud detection
(a) BS+CF+PL+
Summary
(b) BS+CF+PL+
Summary+Text
Earnings forecasting
(c) BS+CF+PL+
Summary
(d) BS+CF+PL+
Summary+Text
Figure 4: Confusion matrix of Claude 3.7 Sonnet on accounting fraud detection and earnings forecasting.
Industry prediction
In the industry prediction task, all models demonstrated predictive performance significantly higher than random guessing. Furthermore, across all models, increasing the input information from Summary to Summary+BS+CF+PL resulted in improved performance. Notably, Claude 3.5 Sonnet achieved state-of-the-art performance with an accuracy score of 0.41 when provided with Summary+BS+CF+PL. The confusion matrix of Claude 3.5 Sonnet for the industry prediction task is shown in Figure 5. We can observe that increasing the input information improves its performance.
These results indicate that industry prediction is a relatively simpler task compared to accounting fraud detection and earnings forecasting, with predictions often possible using only securities reports. This may be attributed to the distinctive characteristics present in the BS, CF, PL, and Summary statements across different industries, such as how assets are primarily composed of “cash and deposits”, “loans”, and “securities” in specific proportions depending on the industry.
(a) Summary
(b) Summary + BS + CF + PL
Figure 5: Confusion matrix of Claude 3.5 Sonnet on industry prediction.
@@ 5 Contamination
Since the securities reports included in EDINET-Bench are publicly available on the Internet, there is a potential for contamination. As a simple check, we conduct company name prediction and year-wise performance analysis.
Company name prediction
To confirm whether the LLM already knew the content of the securities reports used in the evaluation dataset, we measure the performance of a task where the model is given tabular data (BS, CF, PL, Summary) from the test split of the accounting fraud detection dataset and asked to predict the company names. The prompt is shown in Figure 10. For evaluation, we allowed minor company name variations such as including or omitting “Kabushiki Kaisha” by relying on the LLM-as-a-judge approach. For all models, the accuracy is below 0.05, indicating the difficulty for models to associate tabular data with company names.
@@
Accuracy 0.045 0.050 0.045 0.0 0.005 0.0 0.005
Table 8: Performance of each model on company name prediction. BS+CF+PL+Summary is used as input. The comparison between the ground truth company names and the predicted names was conducted using GPT-4o as a judge to account for inconsistencies in name representations.
Year-wise performance analysis
We divide the test split of accounting fraud detection by fiscal start year, and plot the scores for each year. If contamination were present, we would expect higher performance on older data, but as shown in Figure 6, we did not observe the anticipated trend of declining performance on more recent data.
@@
6 Limitations
Our approach to constructing EDINET-Bench has several limitations:
  1. 1.
    Intrinsic difficulty: Among the tasks in our benchmark, the fraud detection and earnings forecasting tasks may be intrinsically challenging with a performance upper bound, as the LLM relies solely on information from a single annual report for its predictions. According to interviews we conducted with auditors, when detecting fraud, auditors first identify suspicious areas based on statistical information from sources like annual reports. They then access confidential internal information and other detailed data required to calculate the statistics presented in the annual report to verify their accuracy. Therefore, future research directions could explore the development of benchmark task designs that enable the model to utilize information beyond the annual report with novel agentic pipelines. Additionally, benchmark evaluations could be designed based on a rubric evaluation, where model responses are evaluated according to multiple criteria tailored to each specific question (Starace et al., 2025; Arora et al., 2025).
  2. 2.
    Mislabeling: When constructing the benchmark dataset for the accounting fraud detection task, we assume that only cases explicitly reported as fraudulent are labeled as such, while all others are considered non-fraudulent. However, there may be undiscovered fraud cases that remain unreported, introducing potential label noise into the dataset. Additionally, our fraud examples are constructed by having the LLM read the contents of the amended reports and determine whether they are related to fraudulent activities. Due to the hallucination problem inherent in LLMs (Huang et al., 2025) and lack of instruction following abilities, there is a risk that some cases may be incorrectly identified as fraudulent. To evaluate the reliability of these classifications, we manually checked the amended reports labeled as fraud cases. We found that most of them involved corrections typically associated with fraudulent behavior, such as fictitious sales. However, we also observed a few cases involving corrections unrelated to financial statements, such as revisions to the number of board members, which do not directly impact the BS, CF, or PL and therefore do not fit our definition of accounting fraud in Appendix 16.
  3. 3.
    Contamination: Our benchmark dataset is constructed using EDINET. Since EDINET’s published documents are accessible on the Internet, some reports in EDINET-Bench may have been included in training datasets of LLMs. In such cases, the models might already have knowledge of the correct labels (Oren et al., 2024). However, since our benchmark can be automatically constructed, we can update our benchmark with future annual reports that are not part of the LLMs’ pre-training data, thereby mitigating contamination concerns. Also see Sec. 5 for our preliminary contamination analysis.
  4. 4.
    Parsing inconsistency: Although our tool, edinet2dataset, is capable of extracting detailed financial information, such as total assets for each fiscal year from TSV files, variations in terminology and formatting across different securities reports can lead to parsing inconsistencies. As a result, LLMs may interpret the absence of parsed data not as a parsing error, but as if the original securities report lacks the information entirely, potentially leading to false positives in accounting fraud detection. Therefore, it is necessary to specify this aspect in the prompt to avoid misjudgment.
@@ 7 Conclusion
In this work, we introduced EDINET-Bench, an open-source Japanese financial benchmark designed to evaluate the performance of LLMs on challenging financial tasks including accounting fraud detection, earnings forecasting, and industry prediction. We leveraged Japan’s Electronic Disclosure for Investors’ NETwork (EDINET) to automatically construct EDINET-Bench and can easily incorporate future reports using our toolkit, edinet2dataset. Our experiments reveal that even state-of-the-art LLMs perform only slightly better than logistic regression in binary classification for fraud detection and earnings forecasting. These results highlight current LLM limitations in complex financial analysis and underscore the need for dedicated research to address these challenges, including novel agentic pipelines. We hope our benchmark and toolkit will accelerate advancements in applying large language models to complex financial analysis tasks.
Ethical statement and broader impact
@@
We constructed EDINET-Bench using only publicly available data from EDINET. By releasing EDINET-Bench and edinet2dataset, we aim to contribute significantly to the development of LLMs in real-world financial domain applications. However, since the data represents real companies, there is a risk of misuse, such as damaging a company’s reputation. This benchmark is intended solely for advancing LLM applications in finance and must not be used to target or harm any real companies included in the dataset. In some cases, we noticed fairness issues such as LLM referencing the reputation of auditors (e.g., “Their financial statements have been audited by a reputable firm ([auditor name])”) as part of the decision-making process, which may exacerbate bias against newer or smaller auditors. This demonstrates that alignment and fairness considerations benefit from being examined within each specific domain (in our case, finance). Furthermore, continued research into and mitigation of these societal issues is vital for the responsible advancement of the community. Finally, there is a possibility that this dataset could be misused to make fraud more difficult for LLMs to detect, and we strongly discourage any such attempts, especially in real-world applications.

### v2

@@ 4 Evaluation
We evaluate the performance of current state-of-the-art LLMs in a zero-shot setting, and compare them to classical baselines trained on the train split.
@@ 4.1 Evaluation Setup
Models.
We evaluate several LLMs in a zero-shot setting. For closed-source models, we include GPT-4o (gpt-4o-2024-11-20), o4-mini (o4-mini-2025-04-16), GPT-5 (gpt-5-2025-08-07), Claude 3.5 Haiku (claude-3-5-haiku-20241022), Claude 3.5 Sonnet (claude-3-5-sonnet-20241022), and Claude 3.7 Sonnet (claude-3-7-sonnet-20250219(achiam2023gpt; claude35sonnet). For open-source models, we evaluate Kimi-K2, DeepSeek-V3, DeepSeek-R1, and Llama 3.3 70B (kimiteam2025kimik2; liu2024deepseek; guo2025deepseek; grattafiori2024llama). In addition, we evaluate a Llama-3.2-1B model fine-tuned on the train split. For classical models, we use Logistic Regression, Random Forest, and XGBoost (cox1958regression; random_forest; xgboost).
Sampling parameters.
For inference, the temperature parameter is set to 0.0 and the maximum token length for generation is set to 4096. For GPT-4o, the random seed is fixed at 0.
Prompt.
We use the system prompt “You are a financial analyst” across all tasks. For binary classification tasks (accounting fraud detection and earnings forecasting), models are prompted to output: (1) a prediction (0 or 1), (2) a probability value (0 to 1), and (3) reasoning based on securities report information. For multi-class industry prediction, models output the predicted industry class and classification reasoning. The fraud detection prompt is shown in Figure 3, with other task prompts provided in Appendix A.
Prompt for accounting fraud detection Please analyze the following information extracted from a Japanese company’s securities report for any signs of fraudulent activities. Please note that some data may be missing and represented as ”-” due to parsing errors. The report has been verified by a certified public accountant, and the numerical values are consistent and correct from a calculation perspective. Therefore, please focus your analysis on non-numerical inconsistencies or logical red flags that could suggest fraud. Respond in the following format: JSON: “‘json { ”reasoning”: ”string”, ”prob”: float (between 0 and 1, probability that the report is fraudulent), ”prediction”: int (0: No fraud, 1: fraud) } “‘ The extracted securities report is as follows:
Figure 3: Prompt for accounting fraud detection.
Input formats.
We extract the following items from the reports for prediction: Balance Sheet (BS), Cash Flow (CF), Profit and Loss (PL), Consolidated Summary (Summary), and Text Block (Text). These items are concatenated, appended to the end of the instruction prompt, and fed into the models in a zero-shot manner. To assess how information quantity affects LLM performance, we tested three configurations: Summary only, BS+CF+PL+Summary, and all items (BS+CF+PL+Summary+Text). Note that BS, CF, PL, and Summary consist of tabular data and do not include company names or exact year information, whereas Text contains unstructured textual information.
Cost and processing time analysis.
@@
Table 6: Performance of each model on each task in EDINET-Bench. For LLMs, the scores represent the mean $\pm$ standard deviation over three runs. Bold indicates the best score. For industry prediction, the score for input formats that include the text section is omitted, as the text section of annual reports often explicitly states the company’s industry, and such cases are indicated as “–”. denotes models that are trained on the train split.
@@
Claude 3.5 Haiku Summary 0.61 $\pm$ 0.01 0.19 $\pm$ 0.02 0.41 $\pm$ 0.01 -0.02 $\pm$ 0.02 0.09 $\pm$ 0.00
Summary+BS+CF+PL 0.60 $\pm$ 0.01 0.18 $\pm$ 0.03 0.45 $\pm$ 0.00 0.00 $\pm$ 0.05 0.13 $\pm$ 0.00
Summary+BS+CF+PL+Text 0.67 $\pm$ 0.00 0.28 $\pm$ 0.02 0.44 $\pm$ 0.01 -0.02 $\pm$ 0.01
Claude 3.5 Sonnet Summary 0.64 $\pm$ 0.01 0.05 $\pm$ 0.03 0.54 $\pm$ 0.01 0.08 $\pm$ 0.01 0.24 $\pm$ 0.01
Summary+BS+CF+PL 0.63 $\pm$ 0.03 0.18 $\pm$ 0.03 0.55 $\pm$ 0.01 0.10 $\pm$ 0.02 0.41 $\pm$ 0.00
Summary+BS+CF+PL+Text 0.73 $\pm$ 0.02 0.32 $\pm$ 0.02 0.52 $\pm$ 0.02 0.08 $\pm$ 0.02
Claude 3.7 Sonnet Summary 0.59 $\pm$ 0.01 0.10 $\pm$ 0.02 0.55 $\pm$ 0.01 0.06 $\pm$ 0.02 0.24 $\pm$ 0.01
Summary+BS+CF+PL 0.58 $\pm$ 0.02 0.09 $\pm$ 0.04 0.58 $\pm$ 0.01 0.13 $\pm$ 0.01 0.39 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.67 $\pm$ 0.01 0.25 $\pm$ 0.02 0.61 $\pm$ 0.01 0.16 $\pm$ 0.02
GPT-4o Summary 0.59 $\pm$ 0.00 0.16 $\pm$ 0.02 0.40 $\pm$ 0.00 -0.17 $\pm$ 0.00 0.14 $\pm$ 0.01
Summary+BS+CF+PL 0.61 $\pm$ 0.01 0.19 $\pm$ 0.02 0.41 $\pm$ 0.00 -0.13 $\pm$ 0.01 0.19 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.69 $\pm$ 0.01 0.29 $\pm$ 0.02 0.41 $\pm$ 0.00 -0.14 $\pm$ 0.00
o4-mini Summary 0.53 $\pm$ 0.00 0.01 $\pm$ 0.08 0.51 $\pm$ 0.00 0.05 $\pm$ 0.02 0.19 $\pm$ 0.01
Summary+BS+CF+PL 0.52 $\pm$ 0.01 0.04 $\pm$ 0.05 0.54 $\pm$ 0.03 0.13 $\pm$ 0.04 0.27 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.61 $\pm$ 0.01 0.10 $\pm$ 0.05 0.58 $\pm$ 0.01 0.15 $\pm$ 0.01
GPT-5 Summary 0.56 $\pm$ 0.00 0.00 $\pm$ 0.00 0.58 $\pm$ 0.00 0.09 $\pm$ 0.02 0.21 $\pm$ 0.01
Summary+BS+CF+PL 0.62 $\pm$ 0.03 0.08 $\pm$ 0.03 0.62 $\pm$ 0.01 0.20 $\pm$ 0.02 0.31 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.67 $\pm$ 0.01 0.07 $\pm$ 0.04 0.65 $\pm$ 0.01 0.21 $\pm$ 0.01
@@
Kimi-K2 Summary 0.58 $\pm$ 0.02 0.10 $\pm$ 0.08 0.46 $\pm$ 0.00 -0.06 $\pm$ 0.00 0.14 $\pm$ 0.01
Summary+BS+CF+PL 0.54 $\pm$ 0.02 0.08 $\pm$ 0.03 0.53 $\pm$ 0.00 0.04 $\pm$ 0.00 0.26 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.59 $\pm$ 0.00 0.11 $\pm$ 0.05 0.55 $\pm$ 0.02 0.07 $\pm$ 0.07
DeepSeek-V3 Summary 0.61 $\pm$ 0.03 0.21 $\pm$ 0.02 0.39 $\pm$ 0.01 -0.14 $\pm$ 0.02 0.12 $\pm$ 0.00
Summary+BS+CF+PL 0.59 $\pm$ 0.02 0.12 $\pm$ 0.04 0.40 $\pm$ 0.01 -0.13 $\pm$ 0.01 0.15 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.55 $\pm$ 0.03 0.07 $\pm$ 0.06 0.39 $\pm$ 0.01 -0.15 $\pm$ 0.02
DeepSeek-R1 Summary 0.54 $\pm$ 0.04 0.01 $\pm$ 0.08 0.43 $\pm$ 0.01 -0.06 $\pm$ 0.01 0.11 $\pm$ 0.02
Summary+BS+CF+PL 0.56 $\pm$ 0.01 0.09 $\pm$ 0.02 0.46 $\pm$ 0.00 -0.01 $\pm$ 0.01 0.16 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.63 $\pm$ 0.01 0.15 $\pm$ 0.04 0.44 $\pm$ 0.01 -0.05 $\pm$ 0.02
Llama 3.3 70B Summary 0.58 $\pm$ 0.01 0.11 $\pm$ 0.01 0.42 $\pm$ 0.01 -0.09 $\pm$ 0.03 0.10 $\pm$ 0.01
Summary+BS+CF+PL 0.59 $\pm$ 0.00 0.11 $\pm$ 0.01 0.41 $\pm$ 0.01 -0.11 $\pm$ 0.02 0.14 $\pm$ 0.01
Summary+BS+CF+PL+Text 0.56 $\pm$ 0.04 0.06 $\pm$ 0.06 0.42 $\pm$ 0.01 -0.04 $\pm$ 0.01
Llama-3.2-1B SFT Summary 0.61 $\pm$ 0.01 0.18 $\pm$ 0.05 0.52 $\pm$ 0.01 0.03 $\pm$ 0.03 0.25 $\pm$ 0.01
Summary+BS+CF+PL 0.66 $\pm$ 0.02 0.17 $\pm$ 0.01 0.52 $\pm$ 0.01 0.04 $\pm$ 0.02 0.26 $\pm$ 0.02
Summary+BS+CF+PL+Text 0.54 $\pm$ 0.02 0.09 $\pm$ 0.05 0.48 $\pm$ 0.00 -0.02 $\pm$ 0.01
@@
Logistic Summary 0.68 0.17 0.56 0.05 0.27
Summary+BS+CF+PL 0.59 0.08 0.51 -0.01 0.33
Summary+BS+CF+PL+Text 0.57 0.06 0.51 0.03
Random Forest Summary 0.75 0.38 0.54 0.09 0.35
Summary+BS+CF+PL 0.75 0.38 0.53 0.01 0.40
Summary+BS+CF+PL+Text 0.73 0.31 0.52 0.03
XGBoost Summary 0.68 0.36 0.55 0.10 0.34
Summary+BS+CF+PL 0.63 0.27 0.53 0.06 0.42
Summary+BS+CF+PL+Text 0.61 0.24 0.53 0.05
@@ 4.2 Results
Table 4.1 shows the performance of each model on each task in EDINET-Bench. In the following, we discuss the results for each task.
Accounting fraud detection.
In the accounting fraud detection task, the random forest achieves the highest performance. All LLMs struggle, with even the state-of-the-art Claude 3.5 Sonnet only marginally outperforming logistic regression in terms of ROC-AUC. Nevertheless, most models benefite substantially from incorporating textual information, suggesting that narrative disclosures provide useful signals for fraud detection. In some cases, model reasoning explicitly referred to auditor reputation (e.g., “Their financial statements have been audited by a reputable firm, Azusa Audit Corporation”) as part of the decision-making process, highlighting a potential bias towards well-known auditing corporations. We further examine the feature importance of the logistic regression model, as presented in Table 7. This reveals that total comprehensive income and total assets are influential factors in non-fraud predictions, suggesting a bias towards predicting larger companies as non-fraud and smaller ones as fraud.
Earnings forecasting.
For the earnings forecasting task, many LLMs struggle to deliver precise predictions. Even GPT-5 achieves only a ROC-AUC of 0.65, aligning with the low forecasting performance reported by xie2024neuripsdb. This limitation likely stems from the challenge of predicting next-year earnings based solely on the current securities reports. Unlike the fraud detection task, incorporating text information do not enhance performance in earnings forecasting, suggesting a weaker reliance on textual factors such as company size and audit firm reputation.
@@
Industry prediction.
In the industry prediction task, all models demonstrate predictive performance significantly higher than random guessing. Furthermore, across all models, expanding the input from Summary to Summary+BS+CF+PL results in improved performance. Notably, Claude 3.5 Sonnet achieves state-of-the-art performance with an accuracy score of 0.41 when provided with Summary+BS+CF+PL. These results indicate that industry prediction is a relatively simpler task compared to accounting fraud detection and earnings forecasting, with predictions often possible using only securities reports. This likely reflects differences in asset composition, such as varying proportions of cash, loans, and securities across industries.
@@ 5 Contamination
Since the annual reports in EDINET-Bench are publicly available online, there is a potential risk of test set contamination (oren2024proving; golchin2024time). Several methods have been proposed to detect test set contamination, including prompting models to reproduce verbatim examples from the test set (golchin2024time) and perplexity-based approaches (mattern-etal-2023-membership; chan2025mlebench). However, generating verbatim examples is not suitable for assessing whether a model has memorized the correspondence between reports and their labels, and perplexity cannot be computed for closed-source models. To evaluate the potential impact of contamination, we instead focus on company name prediction and year-wise performance analysis.
Company name prediction.
To confirm whether the models already knew the content of the securities reports used in the evaluation dataset, we measure the performance of a task where the model is given tabular data (BS, CF, PL, Summary) from the test split of the accounting fraud detection dataset and asked to predict the company names. The prompt is shown in Figure 7. For evaluation, we accepte minor variations in company names, such as the inclusion or omission of terms like ‘Co., Ltd.”, using the LLM-as-a-judge approach (zheng2023llm_as_a_judge). Table 8 shows the result. For all models, the accuracy is 0.05 or lower, indicating the difficulty for models to associate tabular data with company names.
Table 8: Performance of each model on company name prediction. BS+CF+PL+Summary is used as input. The comparison between the ground truth company names and the predicted names is conducted using GPT-4o as a judge to account for inconsistencies in name representations.
@@
Acc. 0.045 0.050 0.045 0.000 0.005 0.000 0.005
Year-wise performance analysis.
To check for potential test set contamination, we analyzed performance by fiscal year on the accounting fraud detection task. Since models may have seen older data during training due to their knowledge cutoff dates, contamination would manifest as higher performance on older years compared to recent years. We split the test set by fiscal start year and plotted yearly performance in Figure 4. The results show no declining trend for recent years, indicating that test set contamination is unlikely to be affecting our evaluation.
@@
Although the above results suggest that contamination has only a limited impact on the experiments, we cannot conclude that the dataset is entirely free of contamination. Achieving a fully contamination-free evaluation would require using annual reports released after the model’s cutoff date. This is possible with our toolkit, which iteratively updates datasets with the latest annual reports.
@@ 6 Conclusion
In this work, we introduced EDINET-Bench, an open-source Japanese financial benchmark designed to evaluate the performance of LLMs on challenging financial tasks including accounting fraud detection, earnings forecasting, and industry prediction. We leveraged Japan’s Electronic Disclosure for Investors’ NETwork (EDINET) to construct EDINET-Bench and can easily incorporate future reports using our toolkit, edinet2dataset. Our experiments reveal that even state-of-the-art LLMs perform only slightly better than logistic regression in binary classification for fraud detection and earnings forecasting. This suggests that simply providing reports to LLMs in a straightforward setting is insufficient. The results underscore the need for benchmark frameworks that better emulate the environments in which financial professionals operate, incorporating richer scaffolding, realistic simulations, and task-specific reasoning support to enable more effective problem solving. We hope our benchmark and toolkit will accelerate advancements in applying large language models to complex financial analysis tasks.
Future work
Building on this study, several promising directions remain for future exploration.
Multilingual dataset.
EDINET-Bench is constructed from Japanese annual reports. Extending the benchmark to multiple languages would allow for cross-lingual evaluation of financial document understanding and better reflect the global nature of financial markets. For instance, the U.S. EDGAR system provides functionality analogous to EDINET, suggesting that a comparable dataset construction pipeline could be applied to build a multilingual extension of the benchmark.
Agentic framework.
In this work, we evaluated LLMs in a simple setting where only a single annual report was provided as input. In practice, however, financial professionals rely on a much broader range of information, such as internal records including receipts and meeting minutes, reports from peer companies, and even real-time economic news. Future research could therefore explore benchmark task designs that enable models to incorporate information beyond the annual report, for example by leveraging agentic capabilities such as web browsing for up-to-date financial news, local file access for company records, or a Python execution environment for quantitative analysis (wang2025openhands; yang2024sweagent). In addition, rather than restricting evaluations to binary classification, benchmarks could adopt rubric-based evaluation, where model outputs are assessed according to multiple criteria tailored to each specific question (starace2025paperbenchevaluatingaisability; arora2025healthbenchevaluatinglargelanguage).
Ethics statement
@@
We constructed EDINET-Bench using publicly available data from EDINET, which contains no private information, ensuring compliance with privacy regulations. By releasing EDINET-Bench and edinet2dataset, we aim to contribute significantly to the development of LLMs in real-world financial domain applications. However, since the data represents real companies, there is a risk of misuse, such as damaging a company’s reputation. This benchmark dataset is intended purely for advancing LLM applications in practical tasks and should never be used to target or harm actual companies. There is also a possibility that this dataset could be misused to make fraud more difficult for LLMs to detect. We strongly discourage any such attempts, especially in real-world applications.
Reproducibility statement
In this work, we constructed a benchmark dataset and evaluate models using it. To ensure reproducibility and facilitate further research, we publicly release both the dataset and the code used for its construction, allowing others to rebuild and directly use the dataset. For the evaluation, we also release the evaluation code so that the reported results can be fully reproduced.